A 2 kg metallic object is heated to 500 c and then dropped into a bucket containing 20 kg of water initially at 20 c . when equilibrium is reached , the temperature of mixture is 70 c . what is the specific heat of metal ? ( neglect the heat capacity of container . )
1
Expert's answer
2019-01-04T10:21:10-0500

According to thermodynamic equilibrium equation:

Q1 = Q2: the heat released is equal to the heat consumed

Cmmm(Tm – Teq) = Cwmw(Teq – Tw)

Cm = (C_w m_w (T_eq-T_w) )/(m_m 〖(T〗_m-T_eq)) = (4186×20×(70-20))/(2×(500-70)) = 4867 (J/kg×K)